# Learning Manipulation Skills through Robot Chain-of-Thought with Sparse Failure Guidance

### v1

@@ Abstract
The acquisition of manipulation skills through language instruction remains an unresolved challenge. Recently, vision-language models have made significant progress in teaching robots these skills. However, their performance is restricted to a narrow range of simple tasks. In this paper, we propose that vision-language models can provide a superior source of rewards for agents. Our method decomposes complex tasks into simpler sub-goals, enabling better task comprehension and avoiding potential failures with sparse failure guidance. Empirical evidence demonstrates that our algorithm consistently outperforms baselines such as CLIP, LIV, and RoboCLIP. Specifically, our algorithm achieves a 5.4×5.4$\times$5.4 × higher average success rate compared to the best baseline, RoboCLIP, across a series of manipulation tasks. It has shown a comprehensive understanding of a wide range of robotic manipulation tasks.
@@ 1 Introduction
Learning manipulation skills from language instruction is a critical challenge in robot learning. Despite advancements (RT1; RT2; RTH), mastering robot manipulation remains an unresolved problem. Recently, vision-language models such as CLIP (CLIP), LIV (LIV), Flamingo (Flamingo), PaLM (PaLM), and GPT-4V have markedly improved robots’ capabilities in acquiring manipulation skills (LIV; RoboFlamingo; PaLME). However, these models are limited to a narrow range of simple tasks, such as closing a drawer. Additionally, CLIP (MetaCLIP) and LIV (LIV) focus on language and visual goal alignment but lack an understanding of the manipulation process. Conversely, RoboCLIP (RoboCLIP) provides a sparse reward signal using S3D (S3D) which lacks spatial reasoning capabilities and task completion awareness.
This study proposes a superior reward design mechanism that leverages pre-trained vision-language models as a source of rewards. Vision language foundation models can provide prior knowledge for robots on how to manipulate objects and reinforce the task completion process. Consequently, our algorithm comprehends a wide range of manipulation tasks.
By analogy, consider how children learn. They break complex tasks into manageable stages, simplifying the task and facilitating its completion. For instance, to open a door, children decompose the task into sub-goals such as approaching the handle, grasping it, pulling it back, and opening the door. During learning, they also draw lessons from potential failures and successes, guided by parental advice and reinforcement.
Our algorithm draws inspiration from child learning paradigms. To better understand the manipulation task process, we use a robot chain-of-thought to decompose tasks into a series of actionable steps, leveraging prior task knowledge from foundation models. Meanwhile, we provide sparse human failure guidance, highlighting potential failures for the agents. The designed reward mechanism yields higher rewards when the robot performs tasks in line with task descriptions while avoiding potential failure cases. To reinforce the task completion process, we collect success experiences through foundation models to strengthen the policy.
In this paper, we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RTFG). Our RTFG introduces an innovative reward feedback mechanism using vision language foundation models to equip robots with a broad spectrum of manipulation skills. We demonstrate that our approach performs well across a broad spectrum of tasks, achieving a 5.4×5.4$\times$5.4 × performance improvement over the best existing baseline. Ablation studies show that each component of our algorithm contributes to the overall performance.
@@ 2.1 Learning Manipulation Skills for Robotics
Acquiring general manipulation abilities can be facilitated through learning from expert demonstrations, using behavioral cloning (BC) (perceiver; CLIPort; octo) and inverse reinforcement learning (IRL) (GAIL; liu_imitation). BC directly maps the state space to the action space to mimic the expert’s actions. However, BC faces several limitations, including compounding errors (DAgger), the copycat problem (copycat), and out-of-distribution sample shifts (fight_fire). In contrast, IRL derives both the reward function and the policy from expert demonstrations. However, IRL struggles with sample efficiency, particularly in real-world applications (jeong2020self; matas2018sim; xu2023roboninja).
Mastering manipulation skills with language inputs directly presents several challenges, such as a lack of reward signals and inefficiency in understanding vision-language contexts. Recent studies have explored integrating foundation models to enhance the understanding of language instructions (SayCan). Our algorithm is designed to learn manipulation skills from language inputs by effectively leveraging pre-trained vision-language models as a source of rewards.
@@ 2.2 Foundation Models for Robotics
Vision-language foundation models equip robot agents with compact pre-trained representations. R3M (R3M) co-trains video and language instructions in a self-supervised manner (TCN), enhancing the robot’s ability to mimic expert actions. PaLM-E (PaLME) integrates visual observations with large language models, facilitating low-level robot control that includes reasoning about the control process. RT-1 (RT1) uses transformers to better mimic expert behaviors. RoboFlamingo (RoboFlamingo) uses multi-step observation inputs to capture a comprehensive view of observational data. RT-2 (RT2) innovates by conceptualizing robot actions as a form of language. Furthermore, RT-H (RTH) introduces an action hierarchy through language motions, leveraging the chain-of-thought reasoning capabilities of large pre-trained language models. However, while capable of high-level task decomposition, prior models struggle with learning low-level action primitives. Our algorithm addresses this by employing reinforcement learning to master low-level action primitives. Additionally, prior models often lack mechanisms for correcting potential failure cases. Our algorithm addresses this by incorporating sparse failure guidance, thereby strengthening the robustness of our method.
Foundation models significantly enhance robot manipulation capabilities by offering valuable reward signals. CLIP (CLIP), for instance, aligns image representations with language embeddings. Consequently, (MetaCLIP) employs CLIP to measure the reward distance between a robot’s current observation and its linguistic goal. LIV (LIV) innovatively combines VIP (VIP) and CLIP to more effectively learn both visual and linguistic goal representations. RoboCLIP (RoboCLIP) adopts the S3D (S3D) representation model to assess the reward for the final observation of a trajectory. Furthermore, RL-VLM-F (RL_VLM_F) directly utilizes GPT-4V to label preferences for agent rollout data.
However, current models still struggle to master a broad spectrum of robot tasks, often succeeding only in simple tasks such as closing a drawer. Vision-language models like CLIP (CLIP) and LIV (LIV) lack an understanding of the manipulation process. While RoboCLIP (RoboCLIP) offers a sparse reward signal that aids in policy learning, it lacks spatial reasoning capabilities and task completion awareness. RL-VLM-F (RL_VLM_F) excels in spatial reasoning for manipulation tasks but is deficient in process understanding capabilities. In contrast, our algorithm not only comprehends detailed language instructions from the robot chain-of-thought but also considers failure scenarios. Additionally, it leverages VLM’s spatial reasoning ability through success experience regularization, offering a more comprehensive approach to mastering robot tasks.
Figure 1: First, the robot receives the language instruction, e.g., "open the door." Subsequently, our algorithm interprets this instruction using the VLM’s chain-of-thought processing, breaking it down into several detailed prompts. The robot also considers some failure guidance. Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills. Successful experiences are recorded to reinforce effective behaviors using VLM.
@@ 3 Background
Chain-of-Thought (CoT) The CoT (CoT) approach is a reasoning method designed to enhance the problem-solving capabilities of large language models. This method involves decomposing a problem into a series of intermediate steps or thoughts that collectively guide towards a solution, mirroring the human approach to complex problem solving. By explicitly detailing these intermediate steps, large language models such as GPT-4V are better equipped to navigate and resolve complex challenges that require advanced understanding and logical deduction.
VideoCLIP VideoCLIP (VideoCLIP) is a multi-modal model that connects visual and textual information, specifically designed to understand videos by associating them with relevant text descriptions. The model is trained on a diverse dataset of videos and associated texts, allowing it to develop a nuanced understanding of the relationship between video segments and language. In particular, VideoCLIP minimizes the sum of two multi-modal contrastive losses:
@@ 4 Method
Figure 2: First, text embeddings are generated using VLM chain-of-thought and human failure guidance. Next, a moving window slides across the agent’s rollout trajectory, calculating the NCE reward based on both video and text embeddings. Finally, VLM evaluates the final observation. If this observation is deemed successful, this trajectory is recorded in the success experience replay buffer for policy regularization.
In section 4.1, we introduce the robot chain-of-thought interpretation of the command and sparse human failure guidance. In section 4.2, we explain how to define the rewards for policy learning with foundation models. Next, in section 4.3, we discuss policy learning with success experience regularization. Finally, section 4.4 outlines the pipeline of our algorithm.
4.1 Robot Chain-of-Thought with Sparse Failure Guidance
Manipulation skills consist of a series of actionable control steps. For instance, opening a door includes sub-tasks such as approaching the handle, grasping it, pulling it back, and opening the door. Decomposing tasks simplifies the manipulation process, affording robots a greater level of understanding and control. However, robots may still encounter potential failures during learning. Offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills.
We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $z_{l}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks.
In our algorithm, $z_{l}^{-}$ is a set of offline potential failure case descriptions consisting of 2 to 4 prompts. To enable the agent to recognize and adjust to potential setbacks, we provide sparse human failure guidance before training the agent.
4.2 Foundation Models as Rewarder
The agent acquires manipulation skills by interacting with an environment, guided by language inputs. A central challenge we address is defining a reward function that effectively reinforces desired behaviors. In this work, we argue that pre-trained vision-language models can provide the agent with prior task knowledge, such as how to manipulate objects. Consequently, we propose a novel reward design mechanism based on robot CoT and sparse failure guidance. Specifically, we explore the usage of VideoCLIP (VideoCLIP) as a vision-language foundation model. VideoCLIP excels in capturing the temporal dynamics of video segments and interpreting associated language prompts, aiming to align robot movements with actions described in the task descriptions.
@@ 4.2 Foundation Models as Rewarder
where $z_{v}$ is the video segment embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task descriptions from robot CoT interpretation, and $z_{l}^{-}$ is a set of potential failure case descriptions guided by humans. The training of VideoCLIP, employing a methodology akin to the InfoNCE loss, aims to distinguish between video segments and corresponding positive or negative texts. We argue that a similar NCE-based reward formulation can leverage the VideoCLIP model rather than the reward definition in RoboCLIP (RoboCLIP). Moreover, the max operation in Equation 3 ensures that robot behaviors aligning with the detailed task descriptions and avoiding failure scenarios yield higher rewards to the agent. Otherwise, the reward signal would penalize undesired behaviors.
4.3 Policy Learning with Success Experience Regularization
Our reward design mechanism motivates the agent to explore the necessary manipulation skills. Although this reward signal performs well in manipulation tasks, the VideoCLIP rewards do not reflect the completion of the manipulation tasks. They only indicate whether the robot’s movements align with detailed language goals. To better reinforce task completion, we introduce policy learning with success experience regularization.
We employ GPT-4V to assess the success of the agent’s final observation in its rollout trajectory. If it deemed successful, this trajectory is archived into a success experience replay buffer, enabling the agent to reinforce this success. Subsequently, we introduce a regularization mechanism aimed at refining the agent’s policy to reinforce successful behavior. The regularization loss is integrated with the policy loss, formulated as follows:
@@ 4.3 Policy Learning with Success Experience Regularization
where $D_{E}$ is the success experience replay buffer and $\lambda$ is the hyper-parameter. By incorporating success experiences into the policy loss calculation, the agent is encouraged to repeat desired behaviors. However, the reward signal for success experiences should also be reinforced to help the agent understand the task completion signal. Therefore, we correct the reward function as follows: $r(s_{T})=r(s_{T})+100$, where $s_{T}$ is the successful final observation of the rollout trajectory.
4.4 Pipeline
Our algorithm begins with GPT-4V interpreting language instructions into detailed task descriptions. Meanwhile, humans guide the agent with several potential failure case descriptions. Subsequently, rewards are calculated following Equation 3 with a moving window that slides over the agent’s rollout trajectories. The agent is then updated with these learned rewards to further develop the skills. Concurrently, GPT-4V evaluates the final observation of the rollout trajectory and assigns a preference label, indicating whether it is a successful observation. If it is successful, we record the trajectory into a success experience replay buffer for policy regularization and correct the reward function.
Figure 2 illustrates our method for computing rewards and success regularization that aid the robot in learning manipulation skills. This involves a moving window that traverses the agent’s rollout trajectory, the computation of NCE rewards using language prompts, and the policy regularization using success experiences gathered through GPT-4V. This comprehensive approach aims to effectively teach the agent desired manipulation skills through reinforcement learning.
@@ 5 Experiment
In this section, we assess the effectiveness of our proposed algorithm by conducting experiments in comparison with several baseline models. The experiments are designed to address two questions: (1) How does our algorithm perform compared to other state-of-the-art algorithms? The results are presented in section 5.1. (2) What are the contributing factors in our algorithm? The ablation studies are presented in section 5.2.
Environments To evaluate the effectiveness of our proposed algorithm, we conduct experiments on a wide array of manipulation tasks using MetaWorld-v2 as the test robotic environment. We select 10 tasks from simple to complex, including drawer, window, faucet, button press, door, and soccer, as described by (ROT). Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
Setup To ensure a fair evaluation across all experiments, we use DrQ-v2 (DrQv2) as the reinforcement learning algorithm. Detailed implementation specifics can be found in the references: CLIP (MetaCLIP), LIV (LIV), and RoboCLIP (RoboCLIP). All experiments, including those involving our proposed algorithm, receive the same language instructions as inputs. Implementation details are provided in the Appendix.
5.1 Overall Comparison
To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing it with established baselines: CLIP (MetaCLIP), LIV (LIV), and RoboCLIP (RoboCLIP). CLIP (as described in (MetaCLIP)) employs CLIP (CLIP) as the vision-language model to measure the distance between current observations and language goals. LIV (LIV) aligns goal representations with language instructions, and RoboCLIP (RoboCLIP) utilizes S3D (S3D) to measure the distance between trajectory videos and language as a form of reward. However, CLIP (CLIP) and LIV (LIV) lack an understanding of the manipulation process, while RoboCLIP (RoboCLIP) lacks spatial reasoning capabilities and task completion awareness. Figure 3 illustrates the training curves, depicting the success rates for each algorithm. Each evaluation involves 20 trajectories, and the experiments were replicated across three different environmental random seeds to introduce variability in the position of objects.
Our findings indicate distinct performance trends across various tasks. RoboCLIP displays promising results in tasks involving drawer and window manipulation, as well as soccer, but it is most effective in the simplest tasks such as drawer closing. CLIP performs well in tasks related to faucet, drawer and window manipulation, and soccer, indicating that CLIP focuses more on task completion rather than the manipulation process. LIV shows competence primarily in tasks involving drawer closure, window opening, and soccer. Our algorithm consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as closing and opening drawers, windows, and doors. These results underscore the effectiveness of our approach in handling diverse and challenging scenarios.

### v2

@@ Abstract
Defining reward functions for skill learning has been a long-standing challenge in robotics. Recently, vision-language models (VLMs) have shown promise in defining reward signals for teaching robots manipulation skills. However, existing works often provide reward guidance that is too coarse, leading to inefficient learning processes. In this paper, we address this issue by implementing more fine-grained reward guidance. We decompose tasks into simpler sub-tasks, using this decomposition to offer more informative reward guidance with VLMs. We also propose a VLM-based self imitation learning process to speed up learning. Empirical evidence demonstrates that our algorithm consistently outperforms baselines such as CLIP, LIV, and RoboCLIP. Specifically, our algorithm achieves a 5.4×5.4$\times$5.4 × higher average success rate compared to the best baseline, RoboCLIP, across a series of manipulation tasks.
Keywords: Manipulation Skills, Vision Language Models, Chain-of-Thought
@@ 1 Introduction
Defining reward functions for learning robotic skills has been a long standing challenge in robotics [1]. Traditional methods typically define the rewards through human knowledge and require significant engineering efforts. Some methods also use expert demonstrations to define rewards [1]. However, this procedure requires expert supervision and can be infeasible in complex scenarios [2] . Recently, the success of vision language models (VLM) has inspired researchers to leverage its power to define rewards for skill learning. For example, RoboCLIP [2] captures the temporal information from robot movements with a video backbone S3D [3].
In spite of the preliminary success of VLM-based reward modeling, a downside of these existing works is that their language guidance is too coarse for learning. Consider the task of opening the door, existing works only use the language embedding of "opening the door" in reward definition to guide the learning process. Such guidance can be too sparse for learning: when the robot is not able to approach the door at the beginning of learning process, the rewarder will always provide a near-zero feedback since the door is not likely to be opened. In contrast, in human world we typically use more fine-grained guidance such as "move towards the door", "move the hand towards the handle" and "grab the handle" during the teaching process. Intuitively, using such decomposition can make the guidance more precise, and will lead to richer and denser feedback.
Based on this intuition, in this paper we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RoboCoT) to address this problem. Our main insight is to decompose the task description into a series of actionable steps and then leverage VLM to define a fine-grained reward to guide skill learning, which can make the learning process more sample efficient and yield higher success rates. We also propose a VLM-based self-imitation procedure to further speed up the learning process. In the experiments, we demonstrate that our approach performs well across a broad spectrum of tasks, achieving a 5.4×5.4$\times$5.4 × performance improvement over the best existing baseline. Ablation studies show that each component of our algorithm contributes to the overall performance.
Figure 1: First, the robot receives the language instruction, e.g., "open the door." Subsequently, our algorithm interprets this instruction using a robot chain-of-thought processing, breaking it down into several detailed prompts. The robot also considers some failure guidance. Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills. Successful experiences are recorded by foundation models to reinforce effective behaviors (i.e., VLM-based self-imitation).
@@ 2.1 Learning Manipulation Skills for Robotics
Acquiring general manipulation abilities can be facilitated through learning from expert demonstrations, using behavioral cloning (BC) [4, 5, 6] or inverse reinforcement learning (IRL) [7, 8]. BC directly maps the state space to the action space to mimic the expert’s actions. However, BC faces several limitations, including compounding errors [9], the copycat problem [10], and out-of-distribution sample shifts [11]. In contrast, IRL derives both the reward function and the policy from expert demonstrations. However, IRL struggles with sample efficiency, particularly in real-world applications [12, 13, 14].
Mastering manipulation skills with language inputs directly presents several challenges, such as a lack of reward signals and inefficiency in understanding vision-language contexts. Recent studies have explored integrating VLMs to enhance the understanding of language instructions [15]. Our algorithm is designed by effectively leveraging VLMs as a source of rewards.
@@ 2.2 Foundation Models for Robotics
VLMs equip robot agents with compact pre-trained representations. R3M [16] co-trains video and language instructions in a self-supervised manner [17], enhancing the robot’s ability to mimic expert actions. PaLM-E [18] integrates visual observations with large language models, facilitating low-level robot control that includes reasoning about the control process. RT-1 [19] uses transformers to better mimic expert behaviors. RoboFlamingo [20] uses multi-step observation inputs to capture a comprehensive view of observational data. RT-2 [21] innovates by conceptualizing robot actions as a form of language. Furthermore, RT-H [22] introduces an action hierarchy through language motions, leveraging the chain-of-thought reasoning capabilities of large language models.
VLMs significantly enhance robot manipulation capabilities by offering valuable reward signals. CLIP [23], for instance, aligns image representations with language embeddings. Consequently, [24] employs CLIP to measure the reward distance between a robot’s current observation and its linguistic goal. LIV [25] innovatively combines VIP [26] and CLIP to more effectively learn both visual and linguistic goal representations. RoboCLIP [2] adopts the S3D [3] representation model to assess the reward for the final observation of a trajectory. Furthermore, RL-VLM-F [27] directly utilizes GPT-4V to label preferences for agent rollout data.
However, current models still struggle to master a broad spectrum of robot tasks, often succeeding only in simple tasks such as closing a drawer. VLMs like CLIP [23] and LIV [25] lack an understanding of the manipulation process. On the other hand, RoboCLIP [2] lacks spatial reasoning capabilities and task completion awareness. RL-VLM-F [27] excels in spatial reasoning for manipulation tasks but is deficient in process understanding capabilities. In contrast, our algorithm not only comprehends detailed language instructions from the robot chain-of-thought but also considers failure scenarios. Additionally, it leverages VLM’s task completion reasoning ability through self imitation process, offering a more comprehensive approach to master robot tasks.
@@ 3 Background
Chain-of-Thought (CoT) The CoT [28] approach is a reasoning method designed to enhance the problem-solving capabilities of large language models. This method involves decomposing a problem into a series of intermediate steps or thoughts that collectively guide towards a solution, mirroring the human approach to complex problem solving. By explicitly detailing these intermediate steps, large language models such as GPT-4V are better equipped to navigate and resolve complex challenges that require advanced understanding and logical deduction.
VideoCLIP VideoCLIP [29] is a multi-modal model that connects visual and textual information, specifically designed to understand videos by associating them with relevant text descriptions. The model is trained on a diverse dataset of videos and associated texts, allowing it to develop a nuanced understanding of the relationship between video segments and language. In particular, VideoCLIP minimizes the sum of two multi-modal contrastive losses:
@@ 4 Method
Figure 2: First, text embeddings are generated using robot chain-of-thought (positive prompts) and human failure guidance (negative prompts). Next, a moving window slides across the agent’s rollout trajectory, calculating the NCE reward based on both video and text embeddings. Finally, VLMs evaluate the final observation. If this observation is deemed successful, this trajectory is recorded in the success replay buffer for self imitation.
In section 4.1, we introduce the robot chain-of-thought interpretation. In section 4.2, we explain how to define the rewards for policy learning with VLMs. Finally, in section 4.3, we introduce VLM-based self imitation procedure via success experience regularization.
4.1 Robot Chain-of-Thought
We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $l_{i}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks.
4.2 VLMs as Rewarder
A central challenge we address is defining a reward function that effectively reinforces desired behaviors. In this work, we argue that VLMs can provide the agent with a much more fine-grained reward guidance. Consequently, we propose a novel reward design mechanism based on robot CoT. Specifically, we explore the usage of VideoCLIP [29] as a VLM. VideoCLIP excels in capturing the temporal dynamics of video segments and interpreting associated language prompts, aiming to align robot movements with actions described in the task descriptions.
@@ 4.2 VLMs as Rewarder
where $z_{v}$ is the video segment embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task description embeddings from robot CoT interpretation $l^{+}$, and $z_{l}^{-}$ is a set of potential failure case description embeddings guided by humans. Here, offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills. To enable the agent to recognize and adjust to potential setbacks, we provide sparse human failure guidance before training with 2 to 4 prompts. The training of VideoCLIP, employing a methodology akin to the InfoNCE loss, aims to distinguish between video segments and corresponding positive or negative texts. We argue that a similar NCE-based reward formulation can leverage the VideoCLIP model rather than the reward definition in RoboCLIP [2]. Moreover, Equation 3 ensures that robot behaviors aligning with the detailed task descriptions and avoiding failure scenarios yield higher rewards to the agent. Otherwise, the reward signal would penalize undesired behaviors.
4.3 VLM-based Self Imitation
Our reward design mechanism motivates the agent to explore the necessary manipulation skills. Empirically, we find that the VideoCLIP rewards don’t reflect the completion of the manipulation tasks well. They focus mostly on whether the robot’s movements align with detailed language goals. To better reinforce task completion and speed up the learning, we introduce VLM-based self imitation procedure.
We employ GPT-4V to assess the success of the agent’s final observation in its rollout trajectory. If it deemed successful, this trajectory is archived into a success replay buffer and we correct the reward function as: $r(s_{T})=r(s_{T})+100$. Here, $s_{T}$ is the successful final observation of the rollout trajectory. Subsequently, we introduce a regularization mechanism aimed at refining the agent’s policy to reinforce successful behavior. This motivates the agent to reinforce the success experience. The regularization loss is integrated with the policy loss, formulated as follows:
@@ 4.3 VLM-based Self Imitation
where $D_{E}$ is the success experience replay buffer and $\lambda$ is the hyper-parameter. By incorporating success experiences into the policy loss calculation, the agent is encouraged to repeat desired behaviors.
@@ 5 Experiment
In this section, we assess the effectiveness of our proposed algorithm RoboCoT. The experiments are designed to address two questions: (1) How does our algorithm perform compared to other state-of-the-art algorithms? The results are presented in section 5.1. (2) What are the contributing factors in our algorithm? The ablation studies are presented in section 5.2.
Environments To evaluate the effectiveness of our proposed algorithm RoboCoT, we conduct experiments on a wide array of manipulation tasks using MetaWorld-v2 as the test robotic environment. We select 10 tasks from simple to complex, including drawer, window, faucet, button press, door, and soccer, as described by [30]. Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
Setup To ensure a fair evaluation across all experiments, we use DrQ-v2 [31] as the reinforcement learning algorithm. Detailed implementation specifics can be found in the references: CLIP [24], LIV [25], and RoboCLIP [2]. All experiments, including those involving our proposed algorithm, receive the same language instructions as inputs. Implementation details are provided in the Appendix 8.
5.1 Overall Comparison
@@
To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing it with established baselines: CLIP [24], LIV [25], and RoboCLIP [2]. CLIP [24] employs CLIP [23] as the vision-language model to measure the distance between current observations and language goals. LIV [25] aligns goal representations with language instructions, and RoboCLIP [2] utilizes S3D [3] to measure the distance between trajectory videos and language as a form of reward. However, CLIP [23] and LIV [25] lack an fine-grained reward guidance, while RoboCLIP [2] lacks task completion awareness. Figure 3 illustrates the training curves, depicting the success rates for each algorithm. Each evaluation involves 20 trajectories, and the experiments were replicated across three different environmental random seeds to introduce variability in the position of objects.
Our findings indicate distinct performance trends across various tasks. RoboCLIP displays promising results in tasks involving drawer and window manipulation, as well as soccer, but it is most effective in the simplest tasks such as drawer closing. CLIP performs well in tasks related to faucet, drawer and window manipulation, and soccer, indicating that CLIP focuses more on task completion rather than the manipulation process. LIV shows competence primarily in tasks involving drawer closure, window opening, and soccer. Our algorithm RoboCoT consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as closing and opening drawers, windows, and doors. These results underscore the effectiveness of our approach in handling diverse and challenging scenarios.